OpenStudy (anonymous):

the ages of edna, ellie, elsa are consecutive intergers . the sum of their ages is 108. what are their ages?

OpenStudy (phi):

if call the middle age m then the 3 ages will be m-1 , m and m+1 add them up and you get 108 in other words m-1 + m + m+1 = 108 can you find m, and then m-1 and m+1 ?

OpenStudy (anonymous):

i dont quite understand this?

OpenStudy (phi):

can you simplify m-1 + m + m+1 = 108 ? using algebra

OpenStudy (anonymous):

i tried and no.

OpenStudy (phi):

you can re-arrange things when you add them (order does not matter) so m-1 + m + m+1 = 108 is the same as m+m+m + 1 + -1 = 108 surely you know what 1+ -1 is ???

OpenStudy (anonymous):

yes i do ... 0

OpenStudy (phi):

so you now have m+m+m+0 = 108 adding 0 on the left side does not change anything, so it is the same as m+m+m= 108 if 3 m's add up to 108, what do you figure 1 m is ?

OpenStudy (phi):

if 3 m's are 108, and each m is the same amount, how much is 1 m ?

OpenStudy (phi):

if 3 sodas cost a total of 6.00 , how much did one soda cost ? (this is the same problem, and you use the same idea to solve your "m" problem)

OpenStudy (anonymous):

3.00

OpenStudy (anonymous):

no no wait im wrong

OpenStudy (anonymous):

2.00

OpenStudy (phi):

I hope you meant 2.00 I would do 6.00/3 which gives 2.00 do the same for 3m = 108

OpenStudy (phi):

btw, m+m+m is the same a 3*m (3 times m)

OpenStudy (phi):

if 3 m's add up to 108, each m must be 1/3 of the total using algebra 3m = 108 3m/3 = 108/3 m = 108/3

OpenStudy (phi):

once you figure out what m is, you can find the other ages. they will be m-1 and m+1

OpenStudy (anonymous):

thanks i got it :)